# Exploring the Evidence-Based Beliefs of LLM-Based Programming Assistants

### v2

@@ Abstract
Recent innovations in artificial intelligence (AI), primarily powered by large language models (LLMs), have transformed how programmers develop and maintain software—leading to new frontiers in software engineering (SE). The advanced capabilities of LLM-based programming assistants to support software development tasks have led to a rise in the adoption of LLMs in SE. However, little is known about the evidenced-based practices, tools and processes verified by research findings, supported and adopted by AI programming assistants. To this end, our work conducts a preliminary evaluation exploring the beliefs of LLM used to support software development tasks. We investigate 17 evidence-based claims posited by empirical SE research across five LLM-based programming assistants. Our findings show that LLM-based programming assistants have ambiguous beliefs regarding research claims and lack credible evidence to support responses. Based on our results, we provide implications for practitioners adopting LLM-based programming assistants in development contexts and shed light on future research directions to enhance the reliability and trustworthiness of LLMs—aiming to increase awareness and adoption of evidence-based SE research findings in practice.
@@
LLM-Based Programming Assistants, Evidence-Based Software Engineering, Empirical SE Research
@@ I Introduction
Recent technological advancements in machine learning have transformed how software is constructed and maintained. For example, artificial intelligence (AI)-powered programming assistants leveraging large language models (LLMs) trained on code are increasingly used to support a wide variety of software engineering (SE) tasks, including designing software [1], fixing programming bugs [2], and generating tests [3]. As such, software engineers are increasingly collaborating with LLMs in development workflows. For instance, studies show 84% of developers are familiar with generative AI tools [4] and 76% are or plan to adopt AI tools in development processes [5]. Moreover, GitHub reports approximately half of developers’ code across programming languages is assisted by the GitHub Copilot LLM-based programming assistant [6].
@@
To support software and its engineering, empirical SE research seeks to understand and enhance software development. These findings aim to promote empirically-based decision making in software development contexts, enhancing software products [7]. Yet, software engineers largely ignore research findings in practice—finding the results of empirical studies irrelevant in industry contexts [8, 9]. For example, research shows automated tools are useful for debugging code and provide practical benefits to software projects [10]. Yet, software engineers often avoid using development tools [11] and ignore notifications from these systems [12] in practice.
@@
AI-based programming assistants are increasingly used to automate development tasks and enhance software development workflows [13]. However, little is known about the perceptions (or “beliefs”) of LLM-based programming assistants regarding empirical research claims. As humans and AI increasingly work together to in SE contexts, understanding LLM perceptions of evidence-based practices is crucial. Software influences a growing extent of modern society, and LLMs are increasingly used to support software development [14]—impacting the behavior and decision-making of software engineers [15]. To this end, we aim to answer the following research questions (RQs):
@@
  • RQ1: What do LLM-based programming assistants believe about empirical SE research claims?
  • RQ2: How do LLM-based programming assistants obtain evidence to support their beliefs?
@@
Prior work explores the beliefs and behaviors of human software engineers with regard to empirical SE claims [16]. We extend this work by conducting a conceptual replication study [17] to observe the evidence-based beliefs of LLM-based programming assistants. We test the same fundamental ideas examined by Devanbu et al. [16], prompting AI-powered systems instead of human developers to investigate beliefs regarding evidence-based SE research findings. We found LLM-based programming assistants mostly agree with evidence-based SE claims, but provide ambiguous answers and lack credible sources of evidence. Based on these emerging results, we provide implications for developers utilizing LLM-driven systems to complete development tasks. We also offer future research directions to enhance development workflows and leverage LLMs to improve the adoption of evidence-based practices in software development.
@@
Motivating Example: Consider the LLM-based programming assistant Devin [^1] https://devin.ai/—recently announced as “the first AI software engineer”. Devin is a novel AI system that works with developers to generate code, fix bugs, add new features, and deploy projects in its own development environment [18]. In three demonstration videos for Devin, the AI software engineer uses print statements to debug errors in code, [^2] See: https://www.youtube.com/watch?v=fjHtjT7GO1c (0:52–1:10); https://www.youtube.com/watch?v=TiXAzn2_Xck (2:12–2:25); and https://www.youtube.com/watch?v=ReE2dFJn_uY&t=50s (0:40–0:49) despite being equipped with its own coding editor and common development tools. While print-statement debugging is widely adopted by programmers in practice [19, 20], it is discouraged by research in favor of automated debugging techniques [21].
@@ II-A Terminology
II-A1 LLM-Based Programming Assistants
LLMs are deep learning models trained on large amounts of data [14]. We define LLM-based programming assistants as chat-based systems powered by LLMs that have the capability to assist in coding and SE tasks. For this project, we examine five LLM-based programming assistants: ChatGPT [22] is a chatbot by OpenAI. We accessed ChatGPT 3.5 through its web interface online; GitHub Copilot [23] is a programming assistant powered by OpenAI’s GPT-4 model. We accessed GitHub Copilot Chat via Visual Studio Code; [^3] https://code.visualstudio.com/, Version 1.87.1 Gemini [24] is powered by Google AI and trained on information from Google services. We utilized the Basic version of Gemini 1 online; Blackbox AI [25] claims to be the “best AI model for code”. [^4] https://www.blackbox.ai/ We used Blackbox AI Code Chat version 2.2.10 accessed online; Claude [26] is an AI assistant from Anthropic that supports code generation. We accessed Claude 3 Haiku online.
@@ II-A3 LLM Beliefs
Agent X believes [sentence] S if and only if S is in X’s belief base, or X could prove S, using the belief base and a logic L, within time T belief base. [28, p. 66]
This definition incorporates the capabilities of modern LLMs—which are trained on information provided in a large belief base, but have the ability to generate new knowledge in a fixed amount of time—as opposed to other approaches (i.e., belief as “degrees of confidence”). Further, this definition of belief regarding AI can impact human reasoning [28]. We focus on LLM beliefs with regard to SE research-based claims.
@@ II-B1 Beliefs in SE
Empirical SE is is a subset of SE research based on observed and measured phenomena [29], in software development contexts [27]. Many studies demonstrate empirical SE findings are not adopted by practitioners (i.e., [11]). Passos et al. found developers rely on personal experiences for beliefs [30]. Kitchenham et al. suggest SE can benefit from evidence-based practices [31], and offer steps to promote empirically-based decisions [7]. Prior work also investigates if developers’ beliefs vary based on personality [32] and impact software quality and productivity [33]. Devanbu et al. show varied beliefs among programmers when assessing empirical SE claims [16], which forms the basis of this study. We extend this work by investigating the beliefs of LLM-based programming assistants—which are increasingly used in conjunction with human programmers in SE contexts.
II-B2 LLMs in SE
There is emerging research interest on LLMs in SE. Zheng et al. outline 123 research papers exploring LLMs across SE tasks, including code generation, vulnerability detection, code analysis, and Q&A interactions [34]. LLMs can support various facets of the software development lifecycle, such as refactoring, requirements, and design [35]. Research also outlines limitations of LLMs. For instance, Wang et al. suggest developers find ChatGPT excels at simple coding tasks, but lacks the ability to support other development activities [36]. Studies also show theses systems can be inaccurate in SE contexts [37]. Our study investigates LLM-based programming assistants regarding their ability to support evidence-based SE research claims.
II-B3 LLMs in Research
Prior work investigates LLMs within research contexts. Auer and colleagues assess the ability of ChatGPT to answer research-based questions from varying domains [38]. In SE contexts, research explores if LLMs can emulate humans in surveys [39], replicate empirical SE research findings [40], and offers guidelines for usage [41]. Prior work uses LLMs to provide evidence-based responses to practical SE-related questions through a GPT-powered chatbot [42]. We seek to understand how LLM-based programming assistants perceive empirical SE research claims.
@@ III-A Data Collection
To observe the beliefs of LLM-based programming assistants, we used prompting—a key feature of conversational LLMs where users submit queries and receive generated natural language responses [43]. Replicating prior work exploring human software engineers’ beliefs, we used zero-shot prompting to prompt each LLM regarding empirical SE research claims once without providing examples or clarification [44].
@@ III-A1 RQ1: Beliefs
We observed generated responses from LLMs to determine their beliefs. To prompt each assistant, the empirical SE research claims (see Table I) were transformed into questions. For example, “Code quality (defect occurrence) depends on which programming language is used” was input as “Does code quality (defect occurrence) depend on which programming language is used?” (C1). We initially prompted AI systems to provide a 5-point Likert response from one (Strongly Disagree) to five (Strongly Agree) for each claim, as done in the original study [16]—however, we found most LLMs would generate arguments for each rating without selecting a level of agreement. Thus, we transitioned to prompting with yes or no questions. This still provided ambiguous responses, which we incorporate in our analysis.
@@ III-A2 RQ2: Evidence
To examine how LLMs evince their beliefs, we analyzed generated responses outlining beliefs (see Section III-A1 and Table I). In addition, we prompted LLMs to understand how they came up with responses to claims. Following each prompt asking about empirical research claims, we prompted “How did you arrive at your answer?” to gain further insight into how LLMs formed the basis of their beliefs.
@@ III-B1 RQ1: Beliefs
We used closed coding to analyze generated beliefs, categorizing responses as “Yes” (Y), “No” (N), “Ambiguous” (A), or “Not Applicable” (N/A)—indicating whether the LLM agreed with, disagreed with, or provided a noncommittal response (i.e., did not make a clear stance) to claims. For example, a Y response for C14 signifies LLM output agreed static analysis tools improve code quality. We measured ambiguous responses since we do not re-prompt LLMs to obtain definitive answers. Two authors performed the closed coding on all 85 responses—providing individual ratings then meeting to discuss and resolve disagreements. We observed “fair agreement” (Cohen’s $\kappa=0.398$[45].
To further investigate the evidence-based beliefs of LLMs, we calculate the rate of correct responses (✓%) to observe alignment with research claims and observe which findings are most upheld by LLMs. We also calculate variance ($\sigma^{2}$) to determine the most controversial beliefs among LLMs, similar to the original study [16]. To calculate the variance, the initial categorizations were reclassified as numerical values (Y $\rightarrow$ 1, A $\rightarrow$ 2, and N $\rightarrow$ 3) [46, Chapter 3.2]. Then, variance was calculated using Equation 1.
@@ III-B2 RQ2: Evidence
To answer RQ2, we inspected the original response to claims to determine if LLMs provided evidence—i.e., resources or data—to support their beliefs. The first author scanned responses for links to and clear mentions of evidence sources. The resulting evidence was categorized into the following types: (1) research papers published in academic venues; (2) opinionated blogs; (3) social media posts; (4) training resources, i.e., tutorials and documentation; (5) posts on Q&A platforms; (6) tech reports for non-academic studies; and (7) LLM-generated code examples.
We also prompted LLMs to provide evidence for responses (see Section III-A2). Two authors used open coding to analyze the prompted responses—independently evaluating output to investigate sources of evidence. After developing initial codes, the researchers segmented and grouped responses into categories based on common patterns and recurring themes. We observed four main sources of evidence: research, training data, LLM reasoning, and previous conversations.
Our study materials are available in a public repository. [^5] https://anonymous.4open.science/r/LLM-beliefs-0F39/README.md
@@ III-C Threats to Validity
We study five proprietary LLM-based programming assistants trained on varying data from different organizations. However, our results may not generalize to all LLMs. We examine limited empirical SE research claims, which were motivated by the original study [16]. To enhance generalizability, future work can explore different LLMs (i.e., open-source models), more advanced models (i.e., GPT-4), and additional SE research claims.
@@
AI systems themselves cannot “believe”—as their output relies on models and training data. Our method to collect LLM beliefs relies on prompting, and we use this term for “technical convenience, and not as an analysis of ‘real belief”’ [28, p. 66]. Further inspection and more in-depth knowledge of models is needed to provide better insights on LLM perceptions of empirical research claims.
@@ IV Results
IV-A RQ1: LLM-Generated Beliefs
Our results on the beliefs of LLM-based programming assistants are presented in Table II. Most responses ($n=55$, 64.7%) agree with the evidence-based claims explored in our study. We saw a limited number of negative responses ($n=9$, 10.6%) disagreeing with claims, and noted 22 ambiguous responses (25.9%). We provide further insights on the beliefs and accuracy of LLMs below:
LLM Beliefs
The most supported evidence-based practices were unit testing (C9), assert statements (C13), static analysis tools (C14), coding standards (C15), and code reviews (C16) enhance code quality—to which all LLM-based programming assistants agreed. In addition, each LLM agreed with the claims that more complex code leads to more defects (C11) and quality varies based on project context (C12). The least supported claims had no positive responses. For instance, all LLMs provided ambiguous responses to statically typed languages having fewer bugs (C6) and stronger code ownership leading to higher quality software (C8). In addition, no LLM agreed that merge commits are buggier than other types of commits (C8). C8 and C2 (fixing bugs is riskier than adding new features), received the most negative responses ($n=3$)—despite research evidence supporting these claims.
The most controversial beliefs are claims with the highest variance, indicating the most disagreement between LLM-based programming assistants. We found C2—fixing defects is more likely to cause defects than adding new features—was the most controversial ($\sigma^{2}=0.96$). While research supports C2 [49], we observed three negative and two positive responses across LLMs. For instance, ChatGPT responded, “fixing defects is often perceived as less risky than adding new features” and Gemini responded, “fixing defects is generally considered less risky because you’re dealing with existing code”. However, positive responses noted “defects are often interconnected, and fixing one defect can have a ripple effect, causing other defects to emerge. This can lead to a cascade of fixes, each with its own set of risks” (Blackbox AI) and “ adding new features typically involves working with a more isolated and well-defined scope, which can be easier to design, implement, and test thoroughly” (Claude).
LLM Accuracy

### v4

# Exploring the Evidence-Based SE Beliefs of Generative AI Tools thanks: 979-8-3315-9147-2/25/\$31.00 ©2025 IEEE

@@ Abstract
Recent innovations in generative artificial intelligence (AI), primarily powered by large language models (LLMs), have transformed how programmers develop and maintain software. The advanced capabilities of generative AI tools to support software development tasks have led to a rise in their adoption within software engineering (SE) workflows. However, little is known about how AI tools perceive evidence-based beliefs and practices supported by research findings. To this end, we conduct a preliminary evaluation conceptually replicating prior work to explore the “beliefs” of generative AI tools used to support software development tasks. We investigate 17 evidence-based claims posited by empirical SE research across five generative AI tools. Our findings show that generative AI tools have ambiguous beliefs regarding research claims and lack credible evidence to support responses. Based on our results, we provide implications for practitioners integrating generative AI-based systems into development contexts and shed light on future research directions to enhance the reliability and trustworthiness of generative AI—aiming to increase awareness and adoption of evidence-based SE research findings in practice.
@@
Generative AI Tools, Evidence-Based Software Engineering, Empirical SE Research
@@ I Introduction
Recent technological advancements have transformed how software is constructed and maintained. For example, generative artificial intelligence (AI) tools, powered by large language models (LLMs), are increasingly capable of supporting software engineering (SE) tasks, including software design [1], repair [2], implementation [3], and testing [4]. As such, developers are increasingly integrating generative AI into SE workflows. For instance, studies show 84% of developers are familiar with generative AI tools [5] and 76% are or plan to adopt them [6]. Moreover, GitHub reports approximately half of developers’ code across programming languages is assisted by GitHub Copilot, an LLM-based programming assistant [7].
@@
To support software and its engineering, empirical SE research seeks to understand and enhance software development. These findings aim to promote empirically-based decision making in SE contexts to enhance software products [8]. Yet, software engineers largely ignore research findings in practice—finding the results to be irrelevant in industry contexts [9, 10]. For example, while research shows automated tools are useful for debugging code and benefiting software projects [11], developers often avoid them [12] and ignore notifications from these systems [13] in practice.
@@
Generative AI is increasingly used to automate tasks within software development workflows [14]. However, little is known about the perceptions (or “beliefs”) of these systems regarding empirical SE research claims. As humans and generative AI increasingly work together in SE contexts, understanding these beliefs is crucial. Software influences a growing extent of modern society, and generative AI is increasingly used to support SE [15]—impacting the behavior and decision-making of software engineers [16]. To this end, we aim to answer the following research questions (RQs):
@@
  • RQ1: What do generative AI tools believe about empirical SE research claims?
  • RQ2: How do generative AI tools obtain evidence to support their beliefs?
@@
Prior work explores the beliefs and behaviors of human software engineers with regard to empirical SE claims [17]. We extend this work by conducting a conceptual replication study [18] to observe the evidence-based beliefs of generative AI tools used in SE tasks. We test the same fundamental ideas examined by Devanbu et al. [17], prompting AI-powered systems instead of human developers to investigate beliefs regarding evidence-based SE research findings. We found generative AI mostly agrees with evidence-based SE claims, but provide ambiguous answers and lack credible sources of evidence. Based on these emerging results, we provide implications for developers utilizing LLM-driven systems to complete development tasks. We also offer research directions to enhance development workflows and leverage generative AI to improve the adoption of evidence-based practices in SE.
@@
Motivating Example: Consider the generative AI-based tool Devin [^1] https://devin.ai/—recently announced as “the first AI software engineer”. Devin is a novel AI system that works with developers to generate code, fix bugs, add new features, and deploy projects in its own development environment [19]. In three demonstration videos for Devin, the AI software engineer uses print statements to debug errors in code, [^2] See: https://www.youtube.com/watch?v=fjHtjT7GO1c (0:52–1:10); https://www.youtube.com/watch?v=TiXAzn2_Xck (2:12–2:25); and https://www.youtube.com/watch?v=ReE2dFJn_uY&t=50s (0:40–0:49) despite being equipped with its own coding editor and common development tools. While print-statement debugging is widely adopted by programmers in practice [20, 21], it is discouraged by research in favor of automated debugging techniques [22].
@@ II-A Terminology
II-A1 Generative AI Tools
LLMs are deep learning models trained on large amounts of data [15]. In this work, generative AI tools are chat-based systems powered by LLMs capable of assisting in coding and SE tasks. For this project, we examine five tools: ChatGPT [23] is a chatbot by OpenAI. We accessed ChatGPT 3.5 through its web interface online; GitHub Copilot [24] is a programming assistant powered by OpenAI’s GPT-4 model. We accessed GitHub Copilot Chat via Visual Studio Code; [^3] https://code.visualstudio.com/, Version 1.87.1 Gemini [25] is powered by Google AI and trained on information from Google services. We utilized the Basic version of Gemini 1 online; Blackbox AI [26] claims to be the “best AI model for code”. [^4] https://www.blackbox.ai/ We used Blackbox AI Code Chat version 2.2.10 accessed online; Claude [27] is an AI system from Anthropic that supports code generation. We accessed Claude 3 Haiku online.
@@ II-A3 Generative AI Beliefs
Agent X believes [sentence] S if and only if S is in X’s belief base, or X could prove S, using the belief base and a logic L, within time T belief base. [29, p. 66]
This definition incorporates the capabilities of modern generative AI tools—which are trained on information provided in a large belief base, but have the ability to generate new knowledge in a fixed amount of time—as opposed to other approaches (i.e., belief as “degrees of confidence”). Further, this definition of belief regarding AI can impact human reasoning [29]. We focus on generative AI beliefs with regard to research-based claims related to software development.
@@ II-B1 Beliefs in SE
Empirical SE is a subset of SE research based on observed and measured phenomena [30], in software development contexts [28]. Many studies demonstrate empirical SE findings are not adopted by practitioners (i.e., [12]). Passos et al. found developers rely on personal experiences for beliefs [31]. Researchers suggest SE can benefit from evidence-based practices [32], and offer steps to promote empirically-based decisions [8, 33]. Prior work also investigates if developers’ beliefs vary based on personality [34] and impact software quality and productivity [35]. Devanbu et al. show varied beliefs among programmers when assessing empirical SE claims [17], which forms the basis of this study. We extend this work by investigating the beliefs of generative AI tools—which are increasingly used in conjunction with human programmers in SE contexts.
II-B2 Generative AI in SE
There is emerging research interest on generative AI in SE. Zheng et al. outline 123 research papers exploring LLMs across SE tasks, including code generation, vulnerability detection, code analysis, and Q&A interactions [36]. Generative AI can support various tasks in the software development lifecycle, such as refactoring, coding, and design [37]. Research also outlines limitations of generative AI in SE. For instance, Wang et al. suggest developers find ChatGPT excels at simple coding tasks, but lacks the ability to support other SE activities [38]. Studies also show theses systems can be inaccurate in SE contexts [39]. We extend these efforts to investigate generative AI tools regarding their ability to support evidence-based SE research claims.
II-B3 Generative AI in Research
Prior work investigates generative AI in research contexts. Auer et al. assess the ability of ChatGPT to answer research-based questions from varying domains [40]. In SE contexts, research explores if LLMs can emulate humans in surveys [41], replicate empirical SE research findings [42], and offers guidelines for usage [43]. Prior work uses generative AI to provide evidence-based responses to practical SE-related questions through a GPT-powered chatbot [44]. We seek to understand how generative AI tools perceive empirical SE research claims.
@@ III-A Data Collection
To observe the beliefs of generative AI tools, we used prompting—a key feature of conversational LLMs where users submit queries and receive generated natural language responses [45]. Replicating prior work exploring human software engineers’ beliefs, we used zero-shot prompting to prompt each tool regarding empirical SE research claims once without providing examples or clarification [46].
@@ III-A1 RQ1: Beliefs
We observed responses from generative AI tools to determine their beliefs. To prompt each tool, the empirical SE research claims (see Table I) were transformed into questions. For example, “Code quality (defect occurrence) depends on which programming language is used” was input as “Does code quality (defect occurrence) depend on which programming language is used?” (C1). We initially devised prompts to provide a 5-point Likert response from one (Strongly Disagree) to five (Strongly Agree) for each claim, as done in the original study [17]—however, we found most AI tools generated arguments for each rating without selecting a level of agreement. Thus, we transitioned to prompting with yes or no questions—limiting our results. This still provided ambiguous responses, which we incorporate in our analysis.
@@ III-A2 RQ2: Evidence
To examine how AI tools evince their beliefs, we analyzed generated responses outlining beliefs (see Section III-A1 and Table I). In addition, we prompted tools to understand how they came up with responses. Following each prompt asking about empirical research claims, we prompted “How did you arrive at your answer?” to gain further insight into how generative AI forms the basis of their beliefs.
@@ III-B1 RQ1: Beliefs
We used closed coding to analyze generated beliefs, categorizing responses as “Yes” (Y), “No” (N), “Ambiguous” (A), or “Not Applicable” (N/A)—indicating whether generative AI tool output agreed with, disagreed with, or provided a noncommittal response (i.e., did not make a clear stance) to claims. For example, a Y response for C14 signifies generative AI agreed static analysis tools improve code quality. We measured ambiguous responses since we do not re-prompt models to obtain definitive answers. Two authors performed the closed coding on all 85 responses—providing individual ratings then meeting to discuss and resolve disagreements. We observed “fair agreement” (Cohen’s $\kappa=0.398$[47].
To further investigate the evidence-based beliefs of generative AI, we calculate the rate of correct responses (✓%) to observe alignment with research claims and observe which findings are most upheld by AI tools. We also calculate variance ($\sigma^{2}$) to determine the most controversial beliefs among AI tools, similar to the original study [17]. To calculate the variance, the initial categorizations were reclassified as numerical values (Y $\rightarrow$ 1, A $\rightarrow$ 2, and N $\rightarrow$ 3) [48, Chapter 3.2]. Then, variance was calculated using Equation 1.
@@ III-B2 RQ2: Evidence
To answer RQ2, we inspected the original response to claims to determine if generative AI tools provided evidence—i.e., resources or data—to support their beliefs. The first author scanned responses for links to and clear mentions of evidence sources. The resulting evidence was categorized into the following types: (1) research papers published in academic venues; (2) opinionated blogs; (3) social media posts; (4) training resources, i.e., tutorials and documentation; (5) posts on Q&A platforms; (6) tech reports for non-academic studies; and (7) AI-generated code examples.
We also prompted AI tools to provide evidence for responses (see Section III-A2). Two authors used open coding to analyze the prompted responses—independently evaluating output to investigate sources of evidence. After developing initial codes, the researchers segmented and grouped responses into categories based on common patterns and recurring themes. We observed four main sources of evidence: research, training data, model reasoning, and previous conversations. Our study materials are available in a public repository. [^5] https://github.com/code-world-no-blanket/LLM-beliefs
@@ III-C Threats to Validity
We study five proprietary generative AI tools trained on varying data from different organizations. However, our results may not generalize to all generative AI tools. We examine limited empirical SE research claims, which were motivated by the original study [17]. To enhance generalizability, future work can explore different LLM types (i.e., open-source models), more advanced generative AI models (i.e., GPT-4), and additional SE research claims.
@@
Generative AI systems themselves cannot “believe”—as their output relies on models and training data. Our method to collect beliefs relies on prompting, and we use this term for “technical convenience, and not as an analysis of ‘real belief”’ [29, p. 66]. Further inspection and more in-depth knowledge of models is needed to provide better insights on AI perceptions of empirical research claims.
@@ IV Results
IV-A RQ1: Generative AI-Generated Beliefs
Our results on the beliefs of generative AI tools are presented in Table II. Most responses ($n=55$, 64.7%) agree with the evidence-based claims explored in our study. We saw a limited number of negative responses ($n=9$, 10.6%) disagreeing with claims, and noted 22 ambiguous responses (25.9%). We provide further insights on the beliefs and accuracy of generative AI tools below:
Generative AI Beliefs
The most supported evidence-based practices were unit testing (C9), assert statements (C13), static analysis tools (C14), coding standards (C15), and code reviews (C16) enhance code quality—to which all generative AI tools agreed. In addition, each AI tool agreed with the claims that more complex code leads to more defects (C11) and quality varies based on project context (C12). The least supported claims had no positive responses. For instance, all generative AI tools provided ambiguous responses to statically typed languages having fewer bugs (C6) and stronger code ownership leading to higher quality software (C8). In addition, no AI tool agreed that merge commits are buggier than other types of commits (C8). C8 and C2 (fixing bugs is riskier than adding new features), received the most negative responses (n=3n=3)—despite research evidence supporting these claims.
The most controversial beliefs are claims with the highest variance, indicating more disagreement between generative AI tools. We found C2—fixing defects is more likely to cause defects than adding new features—was the most controversial ($\sigma^{2}=0.96$). While research supports C2 [51], we observed three negative and two positive responses. For instance, ChatGPT responded, “fixing defects is often perceived as less risky than adding new features” and Gemini responded, “fixing defects is generally considered less risky because you’re dealing with existing code”. Yet, positive responses noted “defects are often interconnected, and fixing one defect can have a ripple effect, causing other defects to emerge. This can lead to a cascade of fixes, each with its own set of risks” (Blackbox AI) and “ adding new features typically involves working with a more isolated and well-defined scope, which can be easier to design, implement, and test thoroughly” (Claude).
Generative AI Accuracy